Solve for x where x is a real number
27^2x = 9^x-3

Answer by stanbon(48575)

Solve for x where x is a real number
27^2x = 9^x-3
------------------
[3^3]^(2x) = [3^2]^(x-3)
----------------
3^(6x) = 3^(2x-6)
----------------
Equate the exponents to get:
6x = 2x-6
4x = -6
x = -3/2
===============
Cheers,
Stan H.